OpenStudy (anonymous):

Find the number halfway between 3/4 and 3/5

OpenStudy (anonymous):

Im not sure medal me

OpenStudy (johnweldon1993):

Well we can add then and divide by 2...this will give us the average \[\large \frac{3}{4} + \frac{3}{5}\] The least common denominator is 20 so \[\large \frac{15}{20} + \frac{12}{20}\] \[\large \frac{15 + 12}{20}\] \[\large \frac{27}{20}\] That would be your final answer...if you wanted to do the mixed number it would be how many times 20 goes into 27..which is 1...and then the remainder over the denominator \[\large 1 \frac{7}{20}\]